# NRGBoost: Energy-Based Generative Boosted Trees

### v1

@@ E.1 Computational Effort
In Figure 4 we show the training times for NRGBoost as well as the other methods for the best model selected by hyperparameter tuning. We do not report the training times for the RFDE method because it is virtually free when compared to the other methods since the splitting process is random and depends only on the input domain. The data itself is only required for computing leaf probabilities which is inexpensive.
@@
Note that the biggest computational cost for training a NRGBoost model is the Gibbs sampling (accounting for roughly 70% of the training time on average). This could potentially be improved by leveraging higher parallelism than what we used in the experiments (16 virtual cores).
@@
We note also that we believe there is still plenty of margin for optimizing the tree-fitting code used for DEF models. As such, the results presented are merely indicative.
@@
The computational effort for fitting a tree-based model should scale linearly both in the number of samples and the number of features and this is the main source of variation for the training times between datasets. However, we note that larger datasets can benefit more from larger models (e.g., with a larger number of leaves) which are also slower to train.
@@
E.2 MCMC Chain Convergence on MNIST
In Figure 5 we show the convergence of a Gibbs sampler sampling from a NRGBoost model. In only a few samples each chain appears to have converged to the data manifold after starting at a random sample from the initial model (a mixture between the product of training marginals and a uniform). Note how consecutive samples are autocorrelated. In particular it can be rare for a chain to switch between two different modes of the distribution (e.g., switching digits) even though a few such transitions can be observed.
Figure 5: Downsampled MNIST samples generated by Gibbs sampling from a NRGBoost model. Each row corresponds to an independent chain initialized with a sample from the initial model $f_{0}$ (first column). Each column represents a consecutive sample from the chain.

### v2

@@ E.6.1 Training
In Figure 6 we report the training times for NRGBoost, as well as the other methods, for the best model selected by hyperparameter tuning. We do not report the training times for the RFDE method because it is virtually free when compared to the other methods given that the splitting process is random and does not depend on the data, only on the input domain.
@@
Note that the biggest computational cost for training an NRGBoost model is the Gibbs sampling, accounting for roughly 70% of the training time on average. This could potentially be improved by leveraging higher parallelism than what we used (16 virtual cores).
@@
We note also that we believe that there is still plenty of margin for optimizing the tree-fitting code used for DEF models. As such, the results presented here are merely indicative.
@@
The computational effort required for fitting a tree-based model should scale linearly with both the number of samples and the number of features. This is the main source of variation for the training times across datasets. However, we note that larger datasets can benefit more from larger models (e.g., with a larger number of leaves) which also take a longer time to train.
@@
E.6.2 Sampling
In Table 14, we report the time required to draw 10,000 independent samples from each model on each dataset. We do not report sampling times for DEF models because our current implementation of this procedure is very suboptimal, taking even longer than sampling from NRGBoost. We believe, however, that a more optimized approach would, at most, be somewhat slower than sampling from ARF models, since the DEF ensembles are larger.
For NRGBoost, independent samples are generated by starting from an independent sample from $q_{0}$ and then running a fixed number, $B$, of Gibbs sampling iterations, taking only the last iterate as a sample. However, this process only produces an approximate sample from the energy model, as it remains biased by the initial sample, which is drawn from an incorrect distribution. The larger $B$ is, compared to the Markov chain’s mixing time, the more independent the final sample is from the initial one and the smaller this bias. This parameter, therefore, trades-off correctness for computational time, which scales linearly with $B$.
In our experiments we always set $B=100$, but this may be excessive for some datasets. We find that $B=30$ already leads to similar ML efficiency results on most datasets, while cutting down sampling times by a factor of three. Note also that the process of drawing samples is embarrassingly parallel. The numbers we report correspond to running 16 chains in parallel on a CPU with 16 virtual cores. If we were to leverage a CPU with higher core count, we could similarly cut down on sampling times. In fact, the process of sampling from an EBM is not dissimilar to the process of sampling from a diffusion model like TabDDPM, but the latter can leverage the massive parallelism afforded by GPUs.
Table 14: Time required to draw 10000 independent samples in seconds. The last columns shows how much faster a method is compared to NRGBoost in the median case. We find that TabDDPM, the only method that realistically competes with NRGBoost in sample quality is, in the median case, 10x faster for sampling.
AB CH PR AD MBNE MNIST CT $\times$ Faster (Median)
TVAE 0.06 0.05 0.04 0.05 0.22 1.33 0.07 984.0
TabDDPM 42.3 7.36 6.82 1.68 1.13 - 9.69 9.67
ARF 0.38 0.22 0.40 2.62 3.84 - 4.47 55.7
NRGBoost 23.1 49.2 54.6 65.0 194.2 622.4 109.9 -